Problem 4: A box of mass Mis released from rest at location A on the track with a trapezoid cross-section as shown below. The elevated region between A and Bis smooth and so is the elevated region between Cand D. The flat region between B and Cis rough; in this region the motion of the box is retarded. The box slides smoothly from A to B with uniform acceleration of magnitude g sina Go, skids from B to C with uniform deceleration of magnitude ao, and then slides smoothly up from C to rest momentarily at D with acceleration of magnitude g sin 0. At point D, the box reverses its motion, ie, it begins to move in the opposite direction. Assume the following the initial location A is at heights ha relative to the flat region of the track, AB makes angle 0o with the horizontal, CD makes angle 6 with the horizontal and the length of the track BC is E. Also assume that the acceleration of the box in each region has a constant magnitude. Answer the questions below. Outline the physical principles employed for each question in a few steps and express all your answers in terms of the given variables (where necessary) (a) What is the total time taken to travel from A to D? (b) What is the speed of the box midway between Cand D during the forward motion? (c) What range of values of ao will guarantee that the box will stop somewhere in the flat region BC during the reverse motion (d) hat is the maximum height, between A and B, reached by the box in the reverse motion (assuming that such a motion is possible)? What is the optimal value that ao must attain for this to occur?
Answer #1

a)

let time for A to B tAB :

AB = dab = ho /Sin\thetao

a = acceleration = g Sin\thetao

Vo = initial velocity = 0 m/s

using the equation

dab = Vo tAB + (0.5) a tAB2

ho /Sin\thetao = 0 + (0.5) (g Sin\thetao) tAB2

tAB = sqrt(2ho /gSin2\thetao )

Vf = final velocity at B

V2f = V2o + 2a dAB

V2f = 0 + 2(g Sin\thetao) (ho /Sin\thetao)

Vf = sqrt(2gho)

from C to D :

Vo = initial velocity at C = sqrt(2gho - 2aol)

Vf = final velocity at D = 0

a = - g Sin\theta

Using the equation

Vf = Vo + a t

0 = sqrt(2gho - 2aol) - g Sin\theta tCD

tCD = sqrt(2gho - 2aol) /(g Sin\theta)

for B to C :

Vo = initial velocity = sqrt(2gho)

a = - ao

d = BC = l

using

Vf2 = Vo2 + 2 a d

Vf2 = (sqrt(2gho))2 - 2 ao l

Vf = sqrt(2gho - 2aol)

using the equation

Vf = Vo + at

sqrt(2gho - 2aol) = sqrt(2gho) - ao tBC

tBC = (sqrt(2gho) - sqrt(2gho - 2aol) ) /ao

tAD = tAB + tBC + tCD

tAD = sqrt(2ho /gSin2\thetao ) + (sqrt(2gho) - sqrt(2gho - 2aol) ) /ao    + sqrt(2gho - 2aol) /(g Sin\theta)

b)

for C to D

Using the equation

Vf = Vo + a t

Vf = sqrt(2gho - 2aol) - g Sin\theta (sqrt(2gho - 2aol) /(2g Sin\theta))

Vf = sqrt(2gho - 2aol)/2

c)

using

Vf2 = Vo2 + 2 a d

02 = (sqrt(2gho))2 - 2 ao l

ao = gho /l